OpenStudy (anonymous):

Okay, How do I know if the Integral, limits are 6 and infinity, of 1/(x+4)^(1/5) converges or diverges?

OpenStudy (amistre64):

if the integral goes off to iinfinity; its diverging

OpenStudy (anonymous):

So do I integrate it first, then plug in infinity for x then subtract after I plig in 6 for the bottom limit?

OpenStudy (amistre64):

integrate it first, yes

OpenStudy (anonymous):

so I get 3/(5(x-3)^(5/3)?

OpenStudy (amistre64):

i get 4/5 ..... -1/5 + 5/5 = 4/5

OpenStudy (anonymous):

oh woops

OpenStudy (amistre64):

\[\int(x+4)^{-1/5}dx=\frac{5}{4}(x+5)^{4/5}\]

OpenStudy (amistre64):

the 6 is useless in this regards since it adds no real value; and at infinity we just get infinity

OpenStudy (anonymous):

Okay, so when I plug in infinity and subtract the six plugged in, what do I get? This is where i get lost usually

OpenStudy (anonymous):

okay, so 1/infinity would be congerging on zero?

OpenStudy (amistre64):

forget the clutter; (infinity)^4/5 i believe is just infinity

OpenStudy (amistre64):

1/large goes to zero, yes

OpenStudy (anonymous):

but its under 1, so it would be getting smaller

OpenStudy (amistre64):

1/100000 of a piece of pie is a very tiny piece

OpenStudy (amistre64):

this isnt under 1

OpenStudy (anonymous):

under 5 sorry

OpenStudy (amistre64):

\[5\sqrt[5]{(\infty+4)^4} \over 4\]

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

so its diverging

OpenStudy (anonymous):

gotcha thanks alot

OpenStudy (amistre64):

:) good luck

OpenStudy (anonymous):

hey